OpenStudy (anonymous):

If the endpoints of the diameter of a circle are (-10,12) and (12,18), what are the coordinates of the center of the circle?

OpenStudy (curry):

the cordinateest are 1,15 dont forgot medal

OpenStudy (anonymous):

Use the midpoint formula: \[midpoint = (\frac{x_{1} + x_{2}}{2}, \frac{y_{1} + y_{2}}{2})\]\[midpoint = \frac{-10 + 12}{2}, \frac{12+18}{2}\]\[= 1,15\]

OpenStudy (owlfred):

curry, c'mon, just answers don't help